3. A subject’s measured volumes and capacities are as follows: a. TLC = 4800 ml b. VC = 4000 ml c. VT = 350 ml d. ERV = 800 ml i. What is his IRV in ml?
Answer #1

ANSWER.

* Calculation of IRV in ml.

      Formula

            VC = TV+IRV+ERV

      IRV = VC - TV - ERV

   Here

     VC = 4000 ml.

         TV = 350 ml.

         ERV = 800 ml.

As per above formula

          IRV = 4000 - 350 - 800

                 = 2850 ml.